# Rapid Dissipative Ground State Preparation at Chemical Transition States

### v1

@@ II.3 Proof Outline and Complexity Analysis
Next, Lemma 2 gives a tail bound on the ground-state hitting time $\tau_{0}$ for the Markov chain with kernel $P$, assuming the ETH-motivated uniform downward drift condition (see Section B.2 for a formal definition). The uniform downward drift condition reflects the ETH expectation that a local jump operator couples a typical excited eigenstate to a dense set of lower-energy eigenstates with broadly distributed matrix elements. Although the number of accessible lower-energy states grows rapidly with energy, the typical matrix elements to each state decrease accordingly, so that the summed probability of downward energy transitions remains essentially constant.
@@
Lemma 1 (Bounding Successive Ground State Overlap; informal).
@@
At first glance, Eq. 19 seems like a very strong condition, since it is a uniform (worst-case-over-$j$) lower bound rather than a typical-case statement. Two features of our setting make it plausible and, importantly, sufficient for our analysis. First, we do not invoke Eq. 19 from an arbitrary initialization: the previous lemma provides a warm start at step $i$ with state mass already concentrated in the low-energy sector relevant to the subsequent cooling stage. Thus the drift assumption is only required locally—throughout the narrow energy window actually visited during the transition-state portion of the path—rather than globally across the full spectrum. Second, in over-parameterized optimization and related dynamics, a provably close warm start often places the iterate in a regime where the effective dynamics becomes well-conditioned and admits monotone progress guarantees; see [40, 22] and references therein for examples where closeness to a target region upgrades qualitative mixing/escape behavior into quantitative, uniform progress bounds.
@@
where $t_{i}$ denotes the number of kernel applications used at grid point $s_{i}$Section III, we provide a concrete quantum procedure that implements a cooling step realizing such a Markov kernel, together with the corresponding quantum circuit cost. Combining that implementation cost with the step count above yields the overall gate complexity cost, Eq. 10.
@@ III Implementation and Resource analysis
Figure 2: (a) Energy error $\Delta E(N_{T})=\left|\mathrm{tr}\left(H(e^{\tau\mathcal{L}})^{N_{T}}[\rho^{(0)}]\right)-E_{0}\right|$ (a.u.) after $N_{T}$ application of the dissipative time step at each successive Hamiltonian along a discretized path, beginning from the ground state of $H(0)$, $\rho^{(0)}=|E_{0}^{(0)}\rangle\langle E_{0}^{(0)}|$, on the H4 potential energy surface (STO-3G) where we take $\tau=0.01$ in these numerical simulations; the final chemical geometry is the transition state (square H4 molecule) corresponding to $s=0.5$ and the 32nd Hamiltonian along the reaction path. Each marker shows the final energy estimate after $N_{T}$ dissipative time steps per Hamiltonian along the reaction path, with $N_{T}\in\{1,5,10,50\}$ and Hamiltonian indices starting at $H(s_{1})$. (b) Corresponding infidelity $1-\langle E^{(i)}_{0}|\rho^{(i)}|E^{(i)}_{0}\rangle$ on a logarithmic scale for the same runs, showing that larger $N_{T}$ improves preparation fidelity across the path where $|E_{0}^{(i)}\rangle$ is the ground state of the $i$-th reaction path Hamiltonian $H(s_{i})$ and $\rho^{(i)}$ is the approximate ground state prepared via dissipative evolution.
Figure 3: H4 potential energy surface (STO-3G) over rectangular geometries parameterized by $(X,Y)$, with the reaction-path segment connecting $s=0$ weakly correlated starting point to the target transition state (TS) point $s=0.5$. The path follows a rectangle distortion of H4 at fixed area $XY=a^{2}$, with atoms at $(\pm X/2,\pm Y/2,0)$ and aspect ratio swept as $\lambda(s)=\lambda_{\min}^{1-s}\lambda_{\max}^{s}$ with $\lambda_{\min}=0.6$, $\lambda_{\max}=1.6$, and $X=a/\sqrt{\lambda}$, $Y=a\sqrt{\lambda}$ for $s\in[0,1]$; the square geometry with side length $a=1.2$ Å occurs at $s=1/2$ and exhibits strong multi-reference character.
A standard mathematical starting point for engineering Lindbladian dynamics that drive a system into low-energy states is the Davies weak-coupling construction [41, 42]. In this framework one obtains a GKLS generator whose stationary state is the Gibbs state of the target Hamiltonian $H$. Modern dissipative ground-state preparation generalizes this approach to tailor jump operators such that the resulting Lindbladian has the ground state as its steady state. A particularly efficient method for doing so was introduced in Ref. [20] which uses filtered jump operators to construct such a Lindbladian.
@@
over a finite set of unfiltered jump operators $a\in\mathcal{A}$. In the chemical setting, there are several choices for jump operators that lead to efficient mixing [29], e.g., $\mathcal{A}=\{a_{i}^{\dagger}a_{j}\}_{i,j=1}^{N_{o}}$}, where $a_{i}^{\dagger}$ ($a_{j}$) are the second quantized fermion creation (annihilation) operators.
@@ III.1 Time support scaling
Theorem 2 identifies and compares two classes of filter functions that each carry their own distinct asymptotic scaling. In both constructions, the filtered jump operator is realized by a time–domain representation (an integral, approximated by a finite sum) in which copies of the bare jump operator are conjugated by real-time evolution under $H$ and weighted by the chosen filter kernel. The decisive implementation parameter is therefore the kernel’s time support $S(\varepsilon_{\mathrm{leak}})$, i.e., the effective time window over which the kernel has non-negligible weight. This window determines the largest evolution time that must be synthesized and, after choosing a discretization step, the number of distinct time samples appearing in the approximation of $e^{\tau\mathcal{L}}$ for a small time step $\tau>0$ [20].
@@
    for some $\alpha>1$ depending only on the Gevrey class [20].
@@
For Gervey filters, the dominant accuracy parameter is $M_{s}$ the number of quadrature points [20]. For Fourier filters, this parameter is the number of Fourier modes $N_{\omega}$; see Appendix F. The values of these parameters are set by the target ground state leakage $\varepsilon_{\mathrm{leak}}$. In practice, the conversion between target leakage $\varepsilon_{\mathrm{leak}}$ and the accuracy parameter value is straightforward in the case of the Fourier filter e.g., $N_{\omega}\sim 1000$ Fourier modes can tackle spectral gaps in the milliHartrees. In contrast, a Gevrey filter’s optimal parameters are more difficult to determine a priori [20].
III.2 Numerical Demonstration
@@
We now provide a numerical demonstration of the dissipative evolution algorithm. We consider a one-dimensional path on the Born–Oppenheimer potential energy surface (PES) of planar H4. To reduce complexity, we consider a path along which the nuclei remain in a rectangular configuration (i.e., the angle between the axes of the pairs is fixed), reducing the dimension of the energy manifold to a two-dimensional surface, with the aspect ratio between the separation of the pairs varied smoothly as a function of the reaction coordinate $s\in[0,1]$; see Fig. 3. The path is chosen to pass through the configuration in which the nuclei sit in a perfect square (that is, the rectangular configuration has equal length sides).
@@
This square geometry of the H4 molecule is non-trivial from both chemical and algorithmic perspectives: in the STO-3G basis near a side length of $\sim 1.2$ Å, the increased symmetry induces near-degeneracies among competing bonding patterns, producing a strongly multi-reference ground state and a compressed low-energy spectrum [43]. Thus, the rectangular-to-square deformation provides a minimal proxy for a reaction coordinate that crosses a strongly correlated “bottleneck” region, while remaining small enough FCI dimension to allow us to classically simulate the performance of our algorithm. Additionally, we ran our Procrustes alignment procedure described by Proposition 1 in order to establish a consistent orbital gauge along the path. The alignment procedure begins at the end of the reaction path, aligning the basis of $H(s_{N_{H}-2})$ to the “products” geometry $H(s_{N_{H}-1}=1)$, then $H(s_{N_{H}-3})$ is aligned with $H(s_{N_{H}-2})$ and so on until the “reactant” geometry is reached $H(s_{0}=0)$.
@@
We employ quadratic, number-conserving jump operators $A_{ij}^{(1)}=c_{i}^{\dagger}c_{j}+c_{j}^{\dagger}c_{i}$ and $A_{ij}^{(2)}=i(c_{i}^{\dagger}c_{j}-c_{j}^{\dagger}c_{i})$ for $i<j$, allowing for spin mixing between, for example, $\alpha$ and $\beta$ orbitals. With this selected dissipative ansatz, we observe the desired ETH-like behavior in the H4 molecule by visualizing the graph $G=(V,E)$ associated with the Markov process induced by dissipative cooling, which lends credibility to the assumption being made in this regime. The vertices of $G$ are the eigenstates in the FCI subspace $V=\{|E_{i}\rangle\}_{i=0}^{D-1}$ and two eigenstates are connected by an edge $\left(|E_{i}\rangle,|E_{j}\rangle\right)\in E$ if the probability of a transition between the two is non-zero $P_{i\to j}>0$. Fig. 1a depicts a visualization of the graph $G$ with the energy-ordered vertices arranged on a circle, where edges are only drawn if they are greater than a constant taken to be $p_{\min}=0.01$. We can observe that H4 is consistent with ETH at the square multi-reference geometry, as it satisfies the uniform drift condition of Eq. 19.

### v2

@@ II.3 Proof Outline and Complexity Analysis
Next, Lemma 2 gives a tail bound on the ground-state hitting time $\tau_{0}$ for the Markov chain with kernel $P$, assuming the ETH-motivated uniform downward drift condition (see Section B.2 for a formal definition). The uniform downward drift condition reflects the ETH expectation that a local jump operator couples a typical excited eigenstate to a dense set of lower-energy eigenstates with broadly distributed matrix elements. In such settings, convergence does not rely on the absence of local minima but rather on local smoothness and conditioning that induce a monotonic descent direction [jacot2018neural]. Although the number of accessible lower-energy states grows rapidly with energy, the typical matrix elements to each state decrease accordingly, so that the summed probability of downward energy transitions remains essentially constant.
@@
Lemma 1 (Bounding Successive Ground-State Overlap; informal).
@@
At first glance, Eq. 19 seems like a very strong condition, since it is a uniform (worst-case-over-$j$) lower bound rather than a typical-case statement. Two features of our setting make it plausible and, importantly, sufficient for our analysis. First, we do not invoke Eq. 19 from an arbitrary initialization: the previous lemma provides a warm start at step $i$ with state mass already concentrated in the low-energy sector relevant to the subsequent cooling stage. Thus the drift assumption is only required locally—throughout the narrow energy window actually visited during the transition-state portion of the path—rather than globally across the full spectrum. Second, in over-parameterized optimization and related dynamics, a provably close warm start often places the iterate in a regime where the effective dynamics becomes well-conditioned and admits monotone progress guarantees; see [jacot2018neural, nguyen2025Guarantees, Chen2025] and references therein for a full discussion of how closeness to a target region upgrades qualitative mixing/escape behavior into quantitative, uniform progress bounds. The downward-drift assumption mirrors the warm-start gradient-descent regime of neural tangent kernel theory, where linearized dynamics induce a contractive flow toward lower loss without requiring the absence of local minima [jacot2018neural, Chen2025].
@@
where $t_{i}$ denotes the number of kernel applications used at grid point $s_{i}$. In Section III, we provide a concrete quantum procedure that implements a cooling step realizing such a Markov kernel, together with the corresponding quantum circuit cost. Combining that implementation cost with the step count above yields the overall gate complexity cost, Eq. 10.
II.4 Reaction Path Optimization
Figure 2: Reaction-path optimization for N3 in the $(\theta,R)$ internal-coordinate plane, demonstrating that our dissipative ground-state transport need not follow a minimum-energy path. Dashed curves show initial guess paths and solid curves show the optimized trajectories obtained by minimizing the path functional $J(\mathbf{R})$, which trades off geometric length, smoothness, and avoidance of conical-intersection (CI) neighborhoods. All paths connect the cyclic $D_{3h}$ endpoint ($\theta\approx 60^{\circ}$, $R\approx 1.30$$\text{\AA }$) to the bent $C_{2v}$ endpoint ($\theta\approx 120^{\circ}$, $R\approx 1.30$$\text{\AA }$). Red $\times$ markers indicate surface-crossing / CI-type structures reported by Kerkines2009, which the optimized paths detour around to improve clearance from small-gap regions.
In quantum chemistry, the term reaction path most commonly refers to a minimum-energy path (MEP) on the potential-energy surface, typically computed by standard path-finding routines such as the nudged elastic band (NEB) method [Sheppard2008MEPReview, HenkelmanJonsson2000NEB, HenkelmanUberuagaJonsson2000CINEB, Jonsson1998NEBChapter]. In contrast, our algorithm does not require the path to be an MEP; we are free to prescribe any continuous path connecting the chosen endpoint geometries. This freedom allows us to optimize the transport path itself, i.e., the curve along which we carry the ground state via dissipative evolution, rather than accepting the MEP by default. The quality of a candidate path is captured by the Davis-Kahan constant $C_{\mathrm{DK}}$, which controls the stability of the ground-state subspace under motion along the path and therefore directly sets the runtime of our algorithm. In order to optimize reaction paths for efficient dissipative evolution, we make use of the Operator-Lipschitz constant, that allows us to utilize the calculus of variations.
Lemma 4 (Chain-rule bound for the Davis-Kahan constant $C_{\mathrm{DK}}$; informal).
Assume $H_{\mathrm{el}}(\mathbf{R})$ is Fréchet differentiable in $\mathbf{R}$ along the image of the path, with operator-Lipschitz constant $L_{\max}$. Then,
$C_{\mathrm{DK}}\leq\frac{L_{\max}}{\Delta_{\min}}\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|\,ds.$ (22)
Proof of Lemma 4 can be found in Appendix J. It follows that we can minimize $C_{\text{DK}}$ by minimizing the length of the reaction path $\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|\,ds$. Lemma 4 motivates a smoothness-regularized reaction path optimization functional,
$\begin{split}J(\mathbf{R}):=&\,\lambda_{\mathrm{len}}\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|ds\\ +&\,\lambda_{\mathrm{curv}}\int_{0}^{1}\|\mathbf{R}^{\prime\prime}(s)\|^{2}ds\\ +&\,\lambda_{\mathrm{CI}}\int_{0}^{1}\Phi_{\mathrm{CI}}(\mathbf{R}(s))\,ds,\end{split}$ (23)
where $\Phi_{\mathrm{CI}}$ is a smooth penalty that is large near a conical intersection (CI) and small away from it. Scalar values $\lambda_{\mathrm{len}},\lambda_{\mathrm{curv}},\lambda_{\mathrm{CI}}>0$ control the importance of reaction path length, curvature and avoidance of conical intersections, respectively. This approach allows us to control how rapidly the ground state changes along the selected reaction path while avoiding chemical geometries for which the spectral gap closes.
The length term is the leading-order control knob for $C_{\mathrm{DK}}$ through Corollary 1. The curvature term suppresses kinks and high-frequency parameterization artifacts, which in practice prevents localized spikes in $\|\mathbf{R}^{\prime}(s)\|$ on the discrete mesh and stabilizes the minimization. The CP term enforces the intended homotopy class by repelling the curve away from the CP neighborhood, which is the regime where $\Delta(s)$ is expected to be smallest.
The N3 molecule provides a concrete example where this reaction-path freedom matters. As shown by the high-level multireference study of Kerkines2009, the low-lying doublet manifold of N3 contains many surface crossings and conical-intersection-type features as the NNN angle bends from the linear region toward cyclic geometries. This implies that a naive path (e.g., a minimum-energy path or a simple linear interpolation in internal coordinates) can easily encounter geometries where the electronic gap to a nearby excited state becomes small, precisely the regime that worsens adiabatic sensitivity and inflates the Davis–Kahan constant [Kerkines2009].
Operationally, we use the catalog of crossing structures and the qualitative location of intersection regions reported for N3 as targets to avoid during optimization. This yields a simple, reproducible workflow: compute or import a coarse map of “problematic” regions (small-gap or intersection neighborhoods), set $\lambda_{\mathrm{CI}}$ to enforce clearance from those regions, and tune $\lambda_{\mathrm{len}},\lambda_{\mathrm{curv}}$ to maintain a short, smooth path.
In Fig. 2, we visualize the effect of optimizing the path functional $J(\mathbf{R})$ on N3. All three candidate paths connect the cyclic $D_{3h}$ region ($\theta\approx 60^{\circ}$, $R\approx 1.30$ Å) to the bent $C_{2v}$ region ($\theta\approx 120^{\circ}$, $R\approx 1.30$ Å), but the optimized trajectories (solid curves) systematically deform away from the intersection “hot spots.” Compared to their initial guesses (dashed curves), the optimized paths remain smooth and avoid sharp turns while making deliberate detours that increase clearance from the labeled crossing regions (notably the cluster near $\theta\sim 106$$112^{\circ}$ and the interior crossings around $\theta\sim 80$$90^{\circ}$). This behavior is consistent with the intended role of the CI penalty term in $J(\mathbf{R})$, trading a modest increase in geometric distance for a more favorable gap profile along the route, thereby reducing adiabatic sensitivity (and hence $C_{\mathrm{DK}}$) for subsequent ground-state transport.
@@ III Implementation and Resource analysis
Figure 3: (a) Energy error $\Delta E(N_{T})=\left|\mathrm{tr}\left(H(e^{\tau\mathcal{L}})^{N_{T}}[\rho^{(0)}]\right)-E_{0}\right|$ (a.u.) after $N_{T}$ application of the dissipative time step at each successive Hamiltonian along a discretized path, beginning from the ground state of $H(0)$, $\rho^{(0)}=|E_{0}^{(0)}\rangle\langle E_{0}^{(0)}|$, on the H4 potential energy surface (STO-3G) where we take $\tau=0.01$ in these numerical simulations; the final chemical geometry is the transition state (square H4 molecule) corresponding to $s=0.5$ and the 32nd Hamiltonian along the reaction path. Each marker shows the final energy estimate after $N_{T}$ dissipative time steps per Hamiltonian along the reaction path, with $N_{T}\in\{1,5,10,35\}$ and Hamiltonian indices starting at $H(s_{1})$. Chem. Acc. indicates where the threshold for chemical accuracy is: $1.6$ mHa. (b) Corresponding infidelity $1-\langle E^{(i)}_{0}|\rho^{(i)}|E^{(i)}_{0}\rangle$ on a logarithmic scale for the same runs, showing that larger $N_{T}$ improves preparation fidelity across the path where $|E_{0}^{(i)}\rangle$ is the ground state of the $i$-th reaction path Hamiltonian $H(s_{i})$ and $\rho^{(i)}$ is the approximate ground state prepared via dissipative evolution.
Figure 4: H4 potential energy surface (STO-3G) over rectangular geometries parameterized by $(X,Y)$, with the reaction-path segment connecting $s=0$ weakly correlated starting point to the target transition state (TS) point $s=0.5$. The path follows a rectangle distortion of H4 at fixed area $XY=a^{2}$, with atoms at $(\pm X/2,\pm Y/2,0)$ and aspect ratio swept as $\lambda(s)=\lambda_{\min}^{1-s}\lambda_{\max}^{s}$ with $\lambda_{\min}=0.6$, $\lambda_{\max}=1.6$, and $X=a/\sqrt{\lambda}$, $Y=a\sqrt{\lambda}$ for $s\in[0,0.5]$; the square geometry with side length $a=1.2$ Å occurs at $s=0.5$ and exhibits strong multi-reference character.
A standard mathematical starting point for engineering Lindbladian dynamics that drive a system into low-energy states is the Davies weak-coupling construction [Davies1974Markovian, Davies1976MarkovianII]. In this framework one obtains a GKLS generator whose stationary state is the Gibbs state of the target Hamiltonian $H$. Modern dissipative ground-state preparation generalizes this approach to tailor jump operators such that the resulting Lindbladian has the ground state as its steady state. A particularly efficient method for doing so was introduced in Ref. [ding2024single] which uses filtered jump operators to construct such a Lindbladian.
@@
over a finite set of unfiltered jump operators $a\in\mathcal{A}$. In the chemical setting, there are several choices for jump operators that lead to efficient mixing [Li_2025], e.g., $\mathcal{A}=\{a_{i}^{\dagger}a_{j}\}_{i,j=1}^{N_{o}}$}, where $a_{i}^{\dagger}$ ($a_{j}$) are the second quantized fermion creation (annihilation) operators.
@@ III.1 Time support scaling
Theorem 2 identifies and compares two classes of filter functions that each carry their own distinct asymptotic scaling. In both constructions, the filtered jump operator is realized by a time–domain representation (an integral, approximated by a finite sum) in which copies of the bare jump operator are conjugated by real-time evolution under $H$ and weighted by the chosen filter kernel. The decisive implementation parameter is therefore the kernel’s time support $S(\varepsilon_{\mathrm{leak}})$, i.e., the effective time window over which the kernel has non-negligible weight. This window determines the largest evolution time that must be synthesized and, after choosing a discretization step, the number of distinct time samples appearing in the approximation of $e^{\tau\mathcal{L}}$ for a small time step $\tau>0$ [ding2024single].
@@
    for some $\alpha>1$ depending only on the Gevrey class [ding2024single].
@@
For Gevrey filters, the dominant accuracy parameter is $M_{s}$ the number of quadrature points [ding2024single]. For Fourier filters, this parameter is the number of Fourier modes $N_{\omega}$; see Appendix F. The values of these parameters are set by the target ground-state leakage $\varepsilon_{\mathrm{leak}}$. In practice, the conversion between target leakage $\varepsilon_{\mathrm{leak}}$ and the accuracy parameter value is straightforward in the case of the Fourier filter e.g., $N_{\omega}\sim 1000$ Fourier modes can tackle spectral gaps in the milliHartrees. In contrast, a Gevrey filter’s optimal parameters are more difficult to determine a priori [ding2024single].
III.2 Numerical Demonstrations
Figure 5: Energy spectra at the cyclobutadiene automerization transition-state geometry (the square C4H4 molecule) for increasing active spaces $\mathrm{CAS}(2,2)$, $\mathrm{CAS}(4,4)$, and $\mathrm{CAS}(6,6)$. In each panel, CASSCF eigenenergies are plotted relative to the ground state, $E_{j}-E_{0}$, as horizontal ticks. Gray dashed guides indicate a uniform discretization into energy “layers” of width $\Delta E=0.01$ (a.u.), used to define the layer index $L(j)=\lfloor(E_{j}-E_{0})/\Delta E\rfloor$ and the total number of layers $L_{\mathrm{tot}}=\lceil(E_{\max}-E_{0})/\Delta E\rceil$. Insets report $L_{\mathrm{tot}}$ and the measured downhill-drift statistic $p_{\min}$ which is independent of system size leading to ETH-like behavior in the chemical system.
H4 Molecule at the Multi-reference Square Geometry
@@ III.2 Numerical Demonstrations
We consider a one-dimensional path on the Born–Oppenheimer potential energy surface (PES) of planar H4. To reduce complexity, we consider a path along which the nuclei remain in a rectangular configuration (i.e., the angle between the axes of the pairs is fixed), reducing the dimension of the energy manifold to a two-dimensional surface, with the aspect ratio between the separation of the pairs varied smoothly as a function of the reaction coordinate $s\in[0,1]$; see Fig. 4. The path is chosen to pass through the configuration in which the nuclei sit in a perfect square (that is, the rectangular configuration has equal length sides).
@@
This square geometry of the H4 molecule is non-trivial from both chemical and algorithmic perspectives: in the STO-3G basis near a side length of $\sim 1.2$ Å, the increased symmetry induces near-degeneracies among competing bonding patterns, producing a strongly multi-reference ground state and a compressed low-energy spectrum [Liu2024-h4]. Thus, the rectangular-to-square deformation provides a minimal proxy for a reaction coordinate that crosses a strongly correlated “bottleneck” region, while remaining small enough FCI dimension to allow us to classically simulate the performance of our algorithm. Additionally, we ran our Procrustes alignment procedure described by Proposition 1 in order to establish a consistent orbital gauge along the path. The alignment procedure begins at the end of the reaction path, aligning the basis of $H(s_{N_{H}-2})$ to the “products” geometry $H(s_{N_{H}-1}=1)$, then $H(s_{N_{H}-3})$ is aligned with $H(s_{N_{H}-2})$ and so on until the “reactant” geometry is reached $H(s_{0}=0)$.
@@
We employ quadratic, number-conserving jump operators $A_{ij}^{(1)}=c_{i}^{\dagger}c_{j}+c_{j}^{\dagger}c_{i}$ and $A_{ij}^{(2)}=i(c_{i}^{\dagger}c_{j}-c_{j}^{\dagger}c_{i})$ for $i<j$, allowing for spin mixing between, for example, $\alpha$ and $\beta$ orbitals. With this selected dissipative ansatz, we observe the desired ETH-like behavior in the H4 molecule by visualizing the graph $G=(V,E)$ associated with the Markov process induced by dissipative cooling, which lends credibility to the assumption being made in this regime. The vertices of $G$ are the eigenstates in the FCI subspace $V=\{|E_{i}\rangle\}_{i=0}^{D-1}$ and two eigenstates are connected by an edge $\left(|E_{i}\rangle,|E_{j}\rangle\right)\in E$ if the probability of a transition between the two is non-zero $P_{i\to j}>0$. Fig. 1a depicts a visualization of the graph $G$ with the energy-ordered vertices arranged on a circle, where edges are only drawn if their transition probability is bound below by $P_{j\to k}\geq 0.01$. We can observe that H4 is consistent with ETH at the square multi-reference geometry, as it satisfies the uniform drift condition of Eq. 19.
@@
Figure 6: Quantum logical resource estimates for a single dissipative time step operator $W(\sqrt{\tau})\approx e^{\tau\mathcal{L}}$ along with the cost of doing QPE across benchmark chemical systems [Low_2025]. The resource estimates are calculated from resource estimates provided in [Low_2025] for QPE and the block encoding $U_{H}$ using the “DFTHC-BLISS-SA” SOTA electronic structure compression method to minimize TT/Toffoli gate counts. Each point corresponds to a row in Table 2, with Toffoli count on the vertical axis and logical qubits on the horizontal axis. Regions of current and projected hardware capabilities are shaded to illustrate which systems and algorithms may be accessible on near- or medium-term fault-tolerant devices [camps2025quantumcomputingtechnologyroadmaps].
ETH-like Behavior in Small Active Spaces
In this work, our foundational posit is that the probability of successfully lowering the energy during the execution of dissipative dynamics is independent of system size; specifically at chemical transition states. Remarkably, we can observe this ETH-like behavior in small chemical systems. Here we consider the cyclobutadiene automerization transition-state geometry, the square C4H4 molecule, in increasing active spaces CAS(2,2), CAS(4,4), and CAS(6,6). In Fig. 5, we show, for a fixed choice of microcanonical window width $\Delta E=0.01$, the distribution of the spectrum over these microcanonical windows. We observe that $p_{\min}$ is independent of system size as we increase the size of the active space. This observation suggests that ETH-like behavior is abundant in the world of chemical systems from nature that satisfy generic symmetries, namely particle conservation.